OpenStudy (anonymous):

simplify 1/2(6x+4)-1/3(3-6x)

jhonyy9 (jhonyy9):

(6x+4) (3-6x) 2(3x-2) 3(1-2x) ------ - -------- = --------- - ---------- = 3x-2 - (1-2x) = 2 3 2 3 = 3x-2 -1 +2x = 5x -3

OpenStudy (anonymous):

could you give some steps i"m having a hard time seeing how you got the answer